Ivanhoe Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $45 throughout the country to loyal alumni of over 3,500 schools. Ivanhoe’s variable costs are 41% of sales; fixed costs are $118,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $13,400 per month. If Ivanhoe were to raise its sales price by 11% to cover these new costs, what would be the new annual breakeven point in sales dollars?
Answer #1

New selling price = 45+11% = 49.95

New variable cost = 45*45% = 20.25

New contribution margin ratio = (49.95-20.25)/49.95 = 59.46%

New fixed cost = (118,000+13,400)*12 = 1,576,800

Break even sales = Fixed cost/Contribution margin ratio

= 1,576,800/59.46%

= 2,651,866

I have rounded some numbers. Comment if you face any issues